Draw the velocity time graph of a car that start with an initial velocity of 4.2ms and accelerate uniformly at 5ms until it attains a maximum velocity of 13.9ms it then maintain the speed for seconds


1
Expert's answer
2021-03-26T11:25:38-0400

The time required for the car to accelerate from 4.2 m/s to 13.9 m/s at 5 m/s2 is


"t=\frac{v_2-v_1}{a}=\frac{13.9-4.2}{5}=1.94\text{ s}."